Henry invested $12,000 in a new restaurant. When the restaurant was sold two years later, he received $27,000. Find his average annual return by solving the equation 12,000(1+r)^2
Answer by checkley77(12844)

12,000(1+r)^2=27,000
(1+r)^2=27,000/12,000
(1+r)^2=2.25
1+r=sqrt2.25
1+r=1.5
5=1.5=1
r=.5 or 50% annual return on the value of the restraunt.
Proof:
12,000(1+.5)^2=27,000
12,000(1.5)^2=27,000
12,000*2.25=27,000
27,000=27,000